# Electromagnetic interactions in elastic neutrino-nucleon scattering

### v2

@@ I Introduction
Nonzero masses and mixing of neutrinos indicate that they have properties and interactions beyond the Standard Model (SM), in which neutrinos can couple to and bosons only. In view of this, searches for neutrino electromagnetic (EM) properties are of particular interest [1, 2, 3, 4]. Indeed, it is known that the neutrino charge radii are predicted to be nonzero even within the SM [5, 6, *Bernabeu02Erratum, 8], while nonzero neutrino magnetic moments arise already in the minimally extended SM [9]. Also, the beyond-SM theories discuss properties of right-handed neutrinos [10]. In the present study, we consider the most general case of Dirac neutrino EM properties: in addition to the charge radius and magnetic moment, we take into account the electric millicharge and anapole and electric dipole moments.
@@
Effects of neutrino EM interactions are among expected manifestations of new physics in the CENS experiments. The data from the CENS experiments, such as COHERENT, CONUS, and Dresden-II, were already used to set limits on the neutrino millicharge, charge radius, and magnetic moment [16, 27, *Cadeddu2018Erratum, 29, 30, 31, 32]. Data from a number of other new experiments aimed at studying CENS are expected to appear in the near future. In order to study neutrino EM properties in CENS experiments, it is necessary to develop a theoretical formalism that would take into account various neutrino and nuclear EM form factors. It is reasonable to begin solving this problem by considering elastic neutrino scattering on the basic constituent of a nucleus, namely a nucleon. Besides, a positively charged nucleon, i.e., a proton, is the simplest nuclear target. It should be also noted in this connection that elastic neutrino-proton scattering is a promising tool for detecting neutrinos from supernovas [33].
@@ IV.1 Neutrino charge radii
![[Uncaptioned image]](x1.png)
![[Uncaptioned image]](x2.png)
![[Uncaptioned image]](x3.png)
@@ IV.2 Neutrino anapole moments
![[Uncaptioned image]](x5.png)
![[Uncaptioned image]](x6.png)
![[Uncaptioned image]](x7.png)
@@ Acknowledgements.
- Giunti and Studenikin [2015] C. Giunti and A. Studenikin, Rev. Mod. Phys. 87, 531 (2015), URL https://link.aps.org/doi/10.1103/RevModPhys.87.531.
- Giunti et al. [2016] C. Giunti, K. A. Kouzakov, Y.-F. Li, A. V. Lokhov, A. I. Studenikin, and S. Zhou, Ann. Phys. 528, 198 (2016), eprint https://onlinelibrary.wiley.com/doi/pdf/10.1002/andp.201500211, URL https://onlinelibrary.wiley.com/doi/abs/10.1002/andp.201500211.
- Studenikin and Kouzakov [2020] A. I. Studenikin and K. A. Kouzakov, Moscow Univ. Phys. Bull. 75, 379 (2020), ISSN 1934-8460, URL https://doi.org/10.3103/S0027134920050215.
- Giunti et al. [2025] C. Giunti, K. Kouzakov, Y. Li, and A. Studenikin, Annu. Rev. Nucl. Part. Sci. 75, 1 (2025), eprint https://doi.org/10.48550/arXiv.2411.03122, URL https://dx.doi.org/10.1088/0954-3899/19/3/005.
- Bernabéu et al. [2000] J. Bernabéu, L. G. Cabral-Rosetti, J. Papavassiliou, and J. Vidal, Phys. Rev. D 62, 113012 (2000), URL https://link.aps.org/doi/10.1103/PhysRevD.62.113012.
- Bernabéu et al. [2002a] J. Bernabéu, J. Papavassiliou, and J. Vidal, Phys. Rev. Lett. 89, 101802 (2002a), URL https://link.aps.org/doi/10.1103/PhysRevLett.89.101802.
- Bernabéu et al. [2002b] J. Bernabéu, J. Papavassiliou, and J. Vidal, Phys. Rev. Lett. 89, 229902(E) (2002b), URL https://link.aps.org/doi/10.1103/PhysRevLett.89.229902.
- Bernabéu et al. [2004] J. Bernabéu, J. Papavassiliou, and J. Vidal, Nucl. Phys. B 680, 450 (2004), ISSN 0550-3213, URL https://www.sciencedirect.com/science/article/pii/S0550321303011003.
- Fujikawa and Shrock [1980] K. Fujikawa and R. E. Shrock, Phys. Rev. Lett. 45, 963 (1980), URL https://link.aps.org/doi/10.1103/PhysRevLett.45.963.
- Drewes [2013] M. Drewes, Int. J. Mod. Phys. E 22, 1330019 (2013), eprint https://doi.org/10.1142/S0218301313300191, URL https://doi.org/10.1142/S0218301313300191.
- Alvarez Ruso et al. [2022] L. Alvarez Ruso, A. M. Ankowski, S. Bacca, A. B. Balantekin, J. Carlson, S. Gardiner, R. Gonzalez-Jimenez, R. Gupta, T. J. Hobbs, M. Hoferichter, et al., in contribution to Snowmass 2021 (2022), eprint 2203.09030.
- Chen [2021] Q. Chen, Ph.D. thesis, Theses and Dissertations–Physics and Astronomy. 86. (2021).
- Tomalak et al. [2021] O. Tomalak, P. Machado, V. Pandey, and R. Plestid, J. High Energy Phys. 2021, 97 (2021), ISSN 1029-8479, URL https://doi.org/10.1007/JHEP02(2021)097.
- Tomalak et al. [2022a] O. Tomalak, Q. Chen, R. J. Hill, and K. S. McFarland, Nat. Commun. 13, 5286 (2022a), ISSN 2041-1723, URL https://doi.org/10.1038/s41467-022-32974-x.
- Tomalak et al. [2022b] O. Tomalak, Q. Chen, R. J. Hill, K. S. McFarland, and C. Wret, Phys. Rev. D 106, 093006 (2022b), URL https://link.aps.org/doi/10.1103/PhysRevD.106.093006.
- Corona et al. [2024] M. A. Corona, M. Cadeddu, N. Cargioli, F. Dordei, and C. Giunti, J. High Energy Phys. 2024, 271 (2024), ISSN 1029-8479, URL https://doi.org/10.1007/JHEP05(2024)271.
- Sufian et al. [2020] R. S. Sufian, K.-F. Liu, and D. G. Richards, J. High Energy Phys. 2020, 136 (2020), ISSN 1029-8479, URL https://doi.org/10.1007/JHEP01(2020)136.
- Megias et al. [2020] G. D. Megias, S. Bolognesi, M. B. Barbaro, and E. Tomasi-Gustafsson, Phys. Rev. C 101, 025501 (2020), URL https://link.aps.org/doi/10.1103/PhysRevC.101.025501.
- Zhang et al. [2020] X. Zhang, T. J. Hobbs, and G. A. Miller, Phys. Rev. D 102, 074026 (2020), URL https://link.aps.org/doi/10.1103/PhysRevD.102.074026.
- Liang and Liu [2020] J. Liang and K.-F. Liu, in Proceedings of 37th International Symposium on Lattice Field Theory — PoS(LATTICE2019) (Sissa Medialab, 2020), LATTICE2019, p. 046, URL http://dx.doi.org/10.22323/1.363.0046.
- Freedman [1974] D. Z. Freedman, Phys. Rev. D 9, 1389 (1974), URL https://link.aps.org/doi/10.1103/PhysRevD.9.1389.
- Kopeliovich and Frankfurt [1974] V. B. Kopeliovich and L. L. Frankfurt, JETP Lett. 19, 145 (1974).
- Akimov et al. [2017] D. Akimov, J. B. Albert, P. An, C. Awe, P. S. Barbeau, B. Becker, V. Belov, A. Brown, A. Bolozdynya, B. Cabrera-Palmer, et al. (COHERENT Collaboration), Science 357, 1123 (2017), eprint https://www.science.org/doi/pdf/10.1126/science.aao0990, URL https://www.science.org/doi/abs/10.1126/science.aao0990.
- Yang et al. [2019] J. Yang, J. A. Hernandez, and J. Piekarewicz, Phys. Rev. C 100, 054301 (2019), URL https://link.aps.org/doi/10.1103/PhysRevC.100.054301.
- Payne et al. [2019] C. G. Payne, S. Bacca, G. Hagen, W. G. Jiang, and T. Papenbrock, Phys. Rev. C 100, 061304(R) (2019), URL https://link.aps.org/doi/10.1103/PhysRevC.100.061304.
- Hoferichter et al. [2020] M. Hoferichter, J. Menéndez, and A. Schwenk, Phys. Rev. D 102, 074018 (2020), URL https://link.aps.org/doi/10.1103/PhysRevD.102.074018.
- Cadeddu et al. [2018] M. Cadeddu, C. Giunti, K. A. Kouzakov, Y. F. Li, A. I. Studenikin, and Y. Y. Zhang, Phys. Rev. D 98, 113010 (2018), URL https://link.aps.org/doi/10.1103/PhysRevD.98.113010.
- Cadeddu et al. [2020a] M. Cadeddu, C. Giunti, K. A. Kouzakov, Y. F. Li, A. I. Studenikin, and Y. Y. Zhang, Phys. Rev. D 101, 059902(E) (2020a), URL https://link.aps.org/doi/10.1103/PhysRevD.101.059902.
- Miranda et al. [2020] O. G. Miranda, D. K. Papoulias, G. S. Garcia, O. Sanders, M. Tórtola, and J. W. F. Valle, J. High Energy Phys. 2020, 130 (2020), ISSN 1029-8479, URL https://doi.org/10.1007/JHEP05(2020)130.
- Cadeddu et al. [2020b] M. Cadeddu, F. Dordei, C. Giunti, Y. F. Li, E. Picciau, and Y. Y. Zhang, Phys. Rev. D 102, 015030 (2020b), URL https://link.aps.org/doi/10.1103/PhysRevD.102.015030.
- Bonet et al. [2022] H. Bonet, A. Bonhomme, C. Buck, K. Fülber, J. Hakenmüller, J. Hempfling, G. Heusser, T. Hugle, M. Lindner, W. Maneschg, et al. (C. O. N. U. S. Collaboration), Eur. Phys. J. C 82, 813 (2022), ISSN 1434-6052, URL https://doi.org/10.1140/epjc/s10052-022-10722-1.
- Atzori Corona et al. [2022] M. Atzori Corona, M. Cadeddu, N. Cargioli, F. Dordei, C. Giunti, Y. F. Li, C. A. Ternes, and Y. Y. Zhang, J. High Energy Phys. 2022, 164 (2022), ISSN 1029-8479, URL https://doi.org/10.1007/JHEP09(2022)164.
- An et al. [2016] F. An, G. An, Q. An, V. Antonelli, E. Baussan, J. Beacom, L. Bezrukov, S. Blyth, R. Brugnera, M. B. Avanzini, et al. (JUNO Collaboration), J. Phys. G: Nucl. Part. Phys. 43, 030401 (2016), URL https://dx.doi.org/10.1088/0954-3899/43/3/030401.
- Popov and Studenikin [2019] A. Popov and A. Studenikin, Eur. Phys. J. C 79, 144 (2019), ISSN 1434-6052, URL https://doi.org/10.1140/epjc/s10052-019-6657-z.
- Popov and Studenikin [2021] A. Popov and A. Studenikin, Phys. Rev. D 103, 115027 (2021), URL https://link.aps.org/doi/10.1103/PhysRevD.103.115027.
- Nowakowski et al. [2005] M. Nowakowski, E. A. Paschos, and J. M. Rodríguez, Eur. J. Phys. 26, 545 (2005), URL https://dx.doi.org/10.1088/0143-0807/26/4/001.
- Navas et al. [2024] S. Navas, C. Amsler, T. Gutsche, C. Hanhart, J. J. Hernández-Rey, C. Lourenço, A. Masoni, M. Mikhasenko, R. E. Mitchell, C. Patrignani, et al. (Particle Data Group Collaboration), Phys. Rev. D 110, 030001 (2024), URL https://link.aps.org/doi/10.1103/PhysRevD.110.030001.
- Ternov [2016a] A. I. Ternov, JETP Letters 104, 75 (2016a), ISSN 1090-6487, URL https://doi.org/10.1134/S0021364016140137.
- Ternov [2016b] A. I. Ternov, Phys. Rev. D 94, 093008 (2016b), URL https://link.aps.org/doi/10.1103/PhysRevD.94.093008.
- Foot et al. [1990] R. Foot, G. C. Joshi, H. Lew, and R. R. Volkas, Mod. Phys. Lett. A 05, 2721 (1990), eprint https://doi.org/10.1142/S0217732390003176, URL https://doi.org/10.1142/S0217732390003176.
- Foot et al. [1993] R. Foot, H. Lew, and R. R. Volkas, J. Phys. G: Nucl. Part. Phys. 19, 361 (1993), URL https://dx.doi.org/10.1088/0954-3899/19/3/005.
- Das et al. [2020] A. Das, D. Ghosh, C. Giunti, and A. Thalapillil, Phys. Rev. D 102, 115009 (2020), URL https://link.aps.org/doi/10.1103/PhysRevD.102.115009.
- Babu and Mohapatra [1990] K. S. Babu and R. N. Mohapatra, Phys. Rev. D 41, 271 (1990), URL https://link.aps.org/doi/10.1103/PhysRevD.41.271.
- Raffelt [1999] G. G. Raffelt, Phys. Rep. 320, 319 (1999), ISSN 0370-1573, URL https://www.sciencedirect.com/science/article/pii/S0370157399000745.
- Haxton and Wieman [2001] W. C. Haxton and C. E. Wieman, Annu. Rev. Nucl. Part. Sci. 51, 261 (2001), ISSN 1545-4134, URL https://www.annualreviews.org/content/journals/10.1146/annurev.nucl.51.101701.132458.
@@
- Giunti and Kim [2007] C. Giunti and C. W. Kim, Fundamentals of Neutrino Physics and Astrophysics (Oxford University Press, 2007), ISBN 9780198508717, URL https://doi.org/10.1093/acprof:oso/9780198508717.001.0001.
- Alberico et al. [2009] W. M. Alberico, S. M. Bilenky, C. Giunti, and K. M. Graczyk, Phys. Rev. C 79, 065204 (2009), URL https://link.aps.org/doi/10.1103/PhysRevC.79.065204.
- Papoulias and Kosmas [2016] D. K. Papoulias and T. S. Kosmas, Adv. High Energy Phys. 2016, 1490860 (2016), eprint https://onlinelibrary.wiley.com/doi/pdf/10.1155/2016/1490860, URL https://onlinelibrary.wiley.com/doi/abs/10.1155/2016/1490860.
- Garvey et al. [1993] G. T. Garvey, W. C. Louis, and D. H. White, Phys. Rev. C 48, 761 (1993), URL https://link.aps.org/doi/10.1103/PhysRevC.48.761.
- Liu et al. [2007] J. Liu, R. D. McKeown, and M. J. Ramsey-Musolf, Phys. Rev. C 76, 025202 (2007), URL https://link.aps.org/doi/10.1103/PhysRevC.76.025202.
- Butkevich [2023] A. V. Butkevich, Phys. Rev. D 107, 073001 (2023), URL https://link.aps.org/doi/10.1103/PhysRevD.107.073001.
- Kouzakov and Studenikin [2017] K. A. Kouzakov and A. I. Studenikin, Phys. Rev. D 95, 055013 (2017), URL https://link.aps.org/doi/10.1103/PhysRevD.95.055013.
- Corona et al. [2023] M. A. Corona, W. M. Bonivento, M. Cadeddu, N. Cargioli, and F. Dordei, Phys. Rev. D 107, 053001 (2023), URL https://link.aps.org/doi/10.1103/PhysRevD.107.053001.
- Aprile et al. [2020] E. Aprile, J. Aalbers, F. Agostini, M. Alfonsi, L. Althueser, F. D. Amaro, V. C. Antochi, E. Angelino, J. R. Angevaare, F. Arneodo, et al. (XENON Collaboration), Phys. Rev. D 102, 072004 (2020), URL https://link.aps.org/doi/10.1103/PhysRevD.102.072004.

### v3

@@ I Introduction
Nonzero masses and mixing of neutrinos indicate that they have properties and interactions beyond the Standard Model (SM), in which neutrinos can couple to and bosons only. In view of this, searches for neutrino electromagnetic (EM) properties are of particular interest [28, 27, 44, 25]. Indeed, it is known that the neutrino charge radii are predicted to be nonzero even within the SM [8, 9, *Bernabeu02Erratum, 10], while nonzero neutrino magnetic moments arise already in the minimally extended SM [23]. Also, the beyond-SM theories discuss properties of right-handed neutrinos [19]. In the present study, we consider the most general case of Dirac neutrino EM properties: in addition to the charge radius and magnetic moment, we take into account the electric millicharge and anapole and electric dipole moments.
@@
Effects of neutrino EM interactions are among expected manifestations of new physics in the CENS experiments. The data from the CENS experiments, such as COHERENT, CONUS, and Dresden-II, were already used to set limits on the neutrino millicharge, charge radius, and magnetic moment [17, 14, *Cadeddu2018Erratum, 36, 13, 11, 6]. Data from a number of other new experiments aimed at studying CENS are expected to appear in the near future. In order to study neutrino EM properties in CENS experiments, it is necessary to develop a theoretical formalism that would take into account various neutrino and nuclear EM form factors. It is reasonable to begin solving this problem by considering elastic neutrino scattering on the basic constituent of a nucleus, namely a nucleon. Besides, a positively charged nucleon, i.e., a proton, is the simplest nuclear target. It should be also noted in this connection that elastic neutrino-proton scattering is a promising tool for detecting neutrinos from supernovas [4].
@@ IV.1 Neutrino charge radii
![[Uncaptioned image]](2412.02169v3/x1.png)
![[Uncaptioned image]](2412.02169v3/x2.png)
![[Uncaptioned image]](2412.02169v3/x3.png)
@@ IV.2 Neutrino anapole moments
![[Uncaptioned image]](2412.02169v3/x5.png)
![[Uncaptioned image]](2412.02169v3/x6.png)
![[Uncaptioned image]](2412.02169v3/x7.png)
@@ Acknowledgements.
- [1] (2017) Observation of coherent elastic neutrino-nucleus scattering. Science 357 (6356), pp. 1123–1126. External Links: Document, Link, https://www.science.org/doi/pdf/10.1126/science.aao0990 Cited by: §I.
- [2] (2009-06) Electromagnetic form factors of the nucleon: new fit and analysis of uncertainties. 79, pp. 065204. External Links: Document, Link Cited by: §II.2, §II.3, §II.3.
- [3] (2022-03) Theoretical tools for neutrino scattering: interplay between lattice QCD, EFTs, nuclear physics, phenomenology, and neutrino event generators. In contribution to Snowmass 2021, External Links: Document, 2203.09030 Cited by: §I.
- [4] (2016-02) Neutrino physics with juno. J. Phys. G: Nucl. Part. Phys.Eur. Phys. J.CPhys. Rev. DEur. J. Phys.JETP LettersPhys. Rev. DPhys. Rev. DPhys. Rev. DPhys. Rep.Annu. Rev. Nucl. Part. Sci.Phys. Rev. CAdv. High Energy Phys.Phys. Rev. CPhys. Rev. CPhys. Rev. DPhys. Rev. DPhys. Rev. DPhys. Rev. Lett.Mod. Phys. Lett. AJ. Phys. G: Nucl. Part. Phys.Annu. Rev. Nucl. Part. Sci. 43 (3), pp. 030401. External Links: Document, Link Cited by: §I.
- [5] (2020-10) Excess electronic recoil events in xenon1t. Phys. Rev. D 102, pp. 072004. External Links: Document, Link Cited by: §IV.3.
- [6] (2022-09-20) Impact of the dresden-ii and coherent neutrino scattering data on neutrino electromagnetic properties and electroweak physics. J. High Energy Phys. 2022 (9), pp. 164. External Links: ISSN 1029-8479, Document, Link Cited by: §I, §IV.1, §IV.3.
- [7] (1990-01) Quantization of electric charge from anomaly constraints and a majorana neutrino. 41, pp. 271–277. External Links: Document, Link Cited by: §II.1.
- [8] (2000-11) Charge radius of the neutrino. Phys. Rev. D 62, pp. 113012. External Links: Document, Link Cited by: §I, §II.1, footnote 2.
- [9] (2002-08) Observability of the neutrino charge radius. Phys. Rev. Lett. 89, pp. 101802. External Links: Document, Link Cited by: §I, §II.1, §II.1, footnote 2.
- [10] (2004) The neutrino charge radius as a physical observable. Nucl. Phys.B 680 (1), pp. 450–478. External Links: ISSN 0550-3213, Document, Link Cited by: §I, §II.1, footnote 2.
- [11] (2022-09-13) First upper limits on neutrino electromagnetic properties from the conus experiment. Eur. Phys. J.C 82 (9), pp. 813. External Links: ISSN 1434-6052, Document, Link Cited by: §I, §IV.3.
- [12] (2023-04) Flux-integrated semiexclusive cross sections for charged-current quasielastic and neutral-current elastic neutrino scattering off and a sterile neutrino oscillation study. 107, pp. 073001. External Links: Document, Link Cited by: §II.3.
- [13] (2020-07) Physics results from the first coherent observation of coherent elastic neutrino-nucleus scattering in argon and their combination with cesium-iodide data. Phys. Rev. D 102, pp. 015030. External Links: Document, Link Cited by: §I, §IV.1, §IV.3.
- [14] (2018-12) Neutrino charge radii from coherent elastic neutrino-nucleus scattering. Phys. Rev. D 98, pp. 113010. External Links: Document, Link Cited by: §I, footnote 2.
- [15] (2021-06) Effective field theory applications: from dark matter to neutrino nucleon scattering. Ph.D. Thesis, Theses and Dissertations–Physics and Astronomy. 86.. External Links: Document Cited by: §I.
- [16] (2023-03) New constraint on neutrino magnetic moment and neutrino millicharge from lux-zeplin dark matter search results. 107, pp. 053001. External Links: Document, Link Cited by: §IV.3.
- [17] (2024-05-23) Momentum dependent flavor radiative corrections to the coherent elastic neutrino-nucleus scattering for the neutrino charge-radius determination. J. High Energy Phys. 2024 (5), pp. 271. External Links: ISSN 1029-8479, Document, Link Cited by: §I, §I.
- [18] (2020-12) Neutrino charge constraints from scattering to the weak gravity conjecture to neutron stars. 102, pp. 115009. External Links: Document, Link Cited by: §II.1.
- [19] (2013) The phenomenology of right handed neutrinos. Int. J. Mod. Phys.E 22 (08), pp. 1330019. External Links: Document, Link, https://doi.org/10.1142/S0218301313300191 Cited by: §I.
- [20] (1993-03) Electric-charge quantization. 19 (3), pp. 361. External Links: Document, Link Cited by: §II.1.
- [21] (1990) CHARGE quantization in the standard model and some of its extensions. 05 (32), pp. 2721–2731. External Links: Document, Link, https://doi.org/10.1142/S0217732390003176 Cited by: §II.1.
- [22] (1974-03) Coherent effects of a weak neutral current. Phys. Rev. D 9, pp. 1389–1392. External Links: Document, Link Cited by: §I.
- [23] (1980-09) Magnetic moment of a massive neutrino and neutrino-spin rotation. Phys. Rev. Lett. 45, pp. 963–966. External Links: Document, Link Cited by: §I, §II.1.
- [24] (1993-08) Determination of proton strange form factors from p elastic scattering. 48, pp. 761–765. External Links: Document, Link Cited by: §II.3.
- [25] (2025) Neutrino electromagnetic properties. 75 (), pp. 1–29. External Links: Document, Link, https://doi.org/10.48550/arXiv.2411.03122 Cited by: §I.
@@
- [26] (2007-03) Fundamentals of Neutrino Physics and Astrophysics. Oxford University Press. External Links: ISBN 9780198508717, Document, Link Cited by: §II.2, §II.2.
- [27] (2016) Electromagnetic neutrinos in laboratory experiments and astrophysics. Ann. Phys. 528 (1-2), pp. 198–215. External Links: Document, Link, https://onlinelibrary.wiley.com/doi/pdf/10.1002/andp.201500211 Cited by: §I.
- [28] (2015-06) Neutrino electromagnetic interactions: a window to new physics. Rev. Mod. Phys. 87, pp. 531–591. External Links: Document, Link Cited by: §I, §II.1, §II.1, §II.1, §II.
- [29] (2001) ATOMIC parity nonconservation and nuclear anapole moments. 51 (Volume 51, 2001), pp. 261–293. External Links: Document, Link, ISSN 1545-4134 Cited by: §II.2.
- [30] (2020-10) Coherent elastic neutrino-nucleus scattering: eft analysis and nuclear responses. Phys. Rev. D 102, pp. 074018. External Links: Document, Link Cited by: §I.
- [31] (1974) Isotopic and chiral structure of neutral current. JETP Lett. 19, pp. 145–147. Cited by: §I.
- [32] (2017-03) Electromagnetic properties of massive neutrinos in low-energy elastic neutrino-electron scattering. 95, pp. 055013. External Links: Document, Link Cited by: §III.
- [33] (2020-08) PDFs and neutrino-nucleon scattering from hadronic tensor. In Proceedings of 37th International Symposium on Lattice Field Theory — PoS(LATTICE2019), LATTICE2019, pp. 046. External Links: Link, Document Cited by: §I.
- [34] (2007-08) Global analysis of nucleon strange form factors at low . 76, pp. 025202. External Links: Document, Link Cited by: §II.3.
- [35] (2020-02) New evaluation of the axial nucleon form factor from electron- and neutrino-scattering data and impact on neutrino-nucleus cross sections. Phys. Rev. C 101, pp. 025501. External Links: Document, Link Cited by: §I.
- [36] (2020-05-26) Implications of the first detection of coherent elastic neutrino-nucleus scattering (cevns) with liquid argon. J. High Energy Phys. 2020 (5), pp. 130. External Links: ISSN 1029-8479, Document, Link Cited by: §I.
- [37] (2024-08) Review of particle physics. Phys. Rev. D 110, pp. 030001. External Links: Document, Link Cited by: §II.1, §II.1, §IV.1, §IV.3.
- [38] (2005-04) All electromagnetic form factors. 26 (4), pp. 545. External Links: Document, Link Cited by: §II.2, §II.
- [39] (2016) Impact of nonstandard interactions on neutrino-nucleon scattering. 2016 (1), pp. 1490860. External Links: Document, Link, https://onlinelibrary.wiley.com/doi/pdf/10.1155/2016/1490860 Cited by: §II.2, §II.3.
- [40] (2019-12) Coherent elastic neutrino-nucleus scattering on from first principles. Phys. Rev. C 100, pp. 061304(R). External Links: Document, Link Cited by: §I.
- [41] (2019-02-19) Neutrino eigenstates and flavour, spin and spin-flavour oscillations in a constant magnetic field. 79 (2), pp. 144. External Links: ISSN 1434-6052, Document, Link Cited by: §I.
- [42] (2021-06) Manifestations of nonzero majorana -violating phases in oscillations of supernova neutrinos. 103, pp. 115027. External Links: Document, Link Cited by: §I.
- [43] (1999) Limits on neutrino electromagnetic properties — an update. 320 (1), pp. 319–327. External Links: ISSN 0370-1573, Document, Link Cited by: §II.1.
- [44] (2020-09-01) The electromagnetic properties of the neutrino as a window to new physics. Moscow Univ. Phys. Bull. 75 (5), pp. 379–397. External Links: ISSN 1934-8460, Document, Link Cited by: §I.
[…]
- [46] (2016-07-01) Neutrino induced magnetic moment and spin precession. 104 (2), pp. 75–81. External Links: ISSN 1090-6487, Document, Link Cited by: §II.1.
- [47] (2016-11) Matter-induced magnetic moment and neutrino helicity rotation in external fields. 94, pp. 093008. External Links: Document, Link Cited by: §II.1.
- [48] (2022-11) Theory of qed radiative corrections to neutrino scattering at accelerator energies. Phys. Rev. D 106, pp. 093006. External Links: Document, Link Cited by: §I.
- [49] (2022-09-08) QED radiative corrections for accelerator neutrinos. Nat. Commun. 13 (1), pp. 5286. External Links: ISSN 2041-1723, Document, Link Cited by: §I.
- [50] (2021-02-11) Flavor-dependent radiative corrections in coherent elastic neutrino-nucleus scattering. J. High Energy Phys. 2021 (2), pp. 97. External Links: ISSN 1029-8479, Document, Link Cited by: §I.
- [51] (2019-11) Electroweak probes of ground state densities. Phys. Rev. C 100, pp. 054301. External Links: Document, Link Cited by: §I.
- [52] (2020-10) Unified model of nucleon elastic form factors and implications for neutrino-oscillation experiments. Phys. Rev. D 102, pp. 074026. External Links: Document, Link Cited by: §I.